# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 5. EBCT, LBCT and UBCT of a function in terms of its DDT
However, $\operatorname{DDT}_{F}(b,c)=2$ and $a\neq b$, which gives that $\operatorname{LBCT}_{F}(a,b,c)=0$. Moreover, $\operatorname{DDT}_{F}(a,b)=10$, and $S(a,b)=\{x_{1}=g,x_{2}=g^{10},x_{3}=g^{13},x_{4}=g^{28},x_{5}=g^{55},x_{1}+a=g^{19},x_{2}+a=g^{46},x_{3}+a=g^{34},x_{4}+a=g^{37},x_{5}+a=0\}$. Then, one can check that $c\not\in W(i,j)$ for all $i\neq j\in\{1,2,3,4\}$, implying that $\operatorname{UBCT}_{F}(a,b,c)=0$.
@@
Let $F$ be a function on $\mathbb{F}_{2^{n}}$. Given any $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, the following statements are equivalent:
@@
On the other hand (permuting the indices from $(a,b,c)$ to $(a,c,d)$),
$\operatorname{LBCT}_{F}(a,c,d)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}(4)\ X+Y=c\\ (5)\ F(X+a)+F(Y+a)=d\\ (6)\ F(X)+F(Y)=d\end{cases}\right\}\right\rvert.$
By Equation (4), $Y=X+c$, and we can rewrite Equations (5) and (6) as
$\begin{cases}F(X+a)+F(X+a+c)=d\\ F(X)+F(X+c)=d.\end{cases}$
@@
Furthermore, (permuting the indices from $(a,b,c)$ to $(c,d,b)$),
$\operatorname{UBCT}_{F}(c,d,b)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}(7)\ F(X+c)+F(Y+c)=b\\ (8)\ F(X)+F(Y)=b\\ (9)\ F(X)+F(X+c)=d\end{cases}\right\}\right\rvert.$
@@
Taking $Y=X+a$, we obtain from Equations (7) and (8), respectively:
$\begin{cases}F(X+c)+F(X+a+c)=b\\ F(X)+F(X+a)=b.\end{cases}$
Note that the second equation of this system is identical to Equation (1). Adding the two equations of this system, we obtain:
$F(X)+F(X+c)=F(X+a)+F(X+a+c),$
which, together with Equation (2), renders $F(X+a)+F(X+a+c)=d$, which is identical with Equation (3).
@@
Corollary 5.4.
Let $F$ be a function on $\mathbb{F}_{2^{n}}$. Given any $a,b,c,d$ with $abcd\neq 0$, let
$A_{k}$ =$=\left\{(a,c,d):\exists\,b,\text{ with }\operatorname{EBCT}_{F}(a,b,c,d)=2k\right\}$
$B_{k}$ =$=\left\{(c,d,b):\exists\,a,\text{ with }\operatorname{EBCT}_{F}(a,b,c,d)=2k\right\},$
where $k\in\{1,2,\cdots,\frac{\delta}{2}\}$. Then, $(a,c,d)\in A_{k}$ if and only if $\operatorname{LBCT}_{F}(a,c,d)=2k$, and $(c,d,b)\in B_{k}$ if and only if $\operatorname{UBCT}_{F}(c,d,b)=2k$.
@@ 6. Consequences of our results for APN functions
If $F$ is an APN function over $\mathbb{F}_{2^{n}}$, then for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F}(c,d)&\mbox{if}~{}a=c\mbox{ and }b=d\\ 0&\mbox{otherwise.}\end{cases}$
$\operatorname{LBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}a=b\\ 0&\mbox{otherwise,}\end{cases}$
and,
$\operatorname{UBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if}~{}c=b\\ 0&\mbox{otherwise,}\end{cases}$
@@
For any $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, $(\operatorname{EBCT}_{F}(a,b,c,d))^{2}\leq\operatorname{LBCT}_{F}(a,c,d)\cdot\operatorname{UBCT}_{F}(c,d,b)$. Furthermore, the equality is met for every $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$ if and only if $F$ is APN.
@@
The inequality follows directly from Corollary 5.4. Suppose that $F$ is APN. Then, by Corollary 6.1, $(\operatorname{EBCT}_{F}(a,b,c,d))^{2}=\operatorname{LBCT}_{F}(a,c,d)\cdot\operatorname{UBCT}_{F}(c,d,b)$. Suppose now that $F$ is not APN. Then, for some $c,d\in\mathbb{F}_{2^{n}}^{*}$, there exist at least four distinct solutions of the equation $F(X+c)+F(X)=d$, which we denote by $z_{1},z_{2},z_{1}+c,z_{2}+c$. Let now $a=z_{1}+z_{2},b=F(z_{1})+F(z_{2})+d$. Then, $\operatorname{EBCT}_{F}(a,b,c,d)=0$, while $\operatorname{LBCT}_{F}(a,c,d)\neq 0$ and $\operatorname{UBCT}_{F}(c,d,b)\neq 0$. ∎
Remark 6.3.
For an APN function $F$, one can compute the explicit entries of all the three tables, $\operatorname{EBCT}_{F}(a,b,c,d)$, $\operatorname{LBCT}_{F}(a,b,c)$ and $\operatorname{UBCT}_{F}(a,b,c)$, for all $a,b,c,d\in\mathbb{F}_{2^{n}}$ using Lemma 2.8 and Corollary 6.1.
Remark 6.4.
For an APN function $F$, Corollary 6.1, Lemma 2.8 and Remark  2.6 imply the result from [3, Theorem 1]:
$\operatorname{FBCT}_{F}(a,b)=\begin{cases}0&\mbox{~{}if~{}}ab\neq 0,a\neq b,\\ 2^{n}&\mbox{~{}otherwise}.\end{cases}$
where $a,b\in\mathbb{F}_{2^{n}}$.
@@ 7. Consequences of our results for some well-known functions
In this section, we will give the EBCT, LBCT and UBCT entries of the Gold, Kasami and Bracken-Leander functions for those parameters such that the functions are differentially 4-uniform permutations; in the case of the Gold function, we also state and prove the EBCT, LBCT and UBCT entries for any parameters. We can also give a very short proof for the EBCT, LBCT and UBCT entries of the inverse function for $n$ even, which is the main result in [12, 16] (for $n$ odd, see the previous section). It may be noted that for those $a,b,c\in\mathbb{F}_{2^{n}}$, for which $abc=0$, the explicit values of LBCT and UBCT can be obtained by using Lemma 2.4 and Lemma 2.8, and, for those $a,b,c,d\in\mathbb{F}_{2^{n}}$, for which $abcd=0$, the explicit values of the EBCT can be obtained by using Lemma 2.8. Thus, throughout this section, we shall determine the LBCT and UBCT entries for all those $a,b,c\in\mathbb{F}_{2^{n}}$, for which $abc\neq 0$ and the EBCT entries for all those $a,b,c,d\in\mathbb{F}_{2^{n}}$, for which $abcd\neq 0$.
For clarity, we first state the results we will use to compute the EBCT, LBCT and UBCT entries of differentially $4$-uniform functions using Theorem 5.1 and Corollary 5.4, which we will need for the functions in this section.
@@
Let $F$ be a differentially $4$-uniform function over $\mathbb{F}_{2^{n}}$. Then, for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F}(c,d)&\mbox{ if }a=c,b=d,\mbox{ or }\\ &\quad\operatorname{DDT}_{F}(c,d)=4,a=z_{1}+z_{2},b=F(z_{1})+F(z_{2}),\mbox{ or }\\ &\quad\operatorname{DDT}_{F}(c,d)=4,a=z_{1}+z_{2}+c,b=F(z_{2})+F(z_{2}+c)\\ 0&\mbox{otherwise},\end{cases}$
where $z_{1},z_{1}+c,z_{2},z_{2}+c$ are the four solutions of the equation $F(X+c)+F(X)=d$ if $\operatorname{DDT}_{F}(c,d)=4$. Next,
$\operatorname{LBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}\operatorname{DDT}_{F}(b,c)=2\mbox{ and }a=b,\mbox{ or }\\ &\quad\operatorname{DDT}_{F}(b,c)=4\mbox{ and }a\in\{b,y_{1}+y_{2},y_{1}+y_{2}+b\},\\ 0&\mbox{otherwise},\end{cases}$
where $y_{1},y_{1}+b,y_{2},y_{2}+b$ are the four solutions of the equation $F(X+b)+F(X)=c$ if $\operatorname{DDT}_{F}(b,c)=4$. Further,
$\operatorname{UBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if}~{}\operatorname{DDT}_{F}(a,b)=2\mbox{ and }c=b,\mbox{ or }\\ &\operatorname{DDT}_{F}(a,b)=4\mbox{ and }c\in\{b,F(x_{1})+F(x_{2}),F(x_{1})+F(x_{2})+b\},\\ 0&\mbox{otherwise},\end{cases}$
where $x_{1},x_{1}+a,x_{2},x_{2}+a$ are the four solutions of the equation $F(X+a)+F(X)=b$ if $\operatorname{DDT}_{F}(a,b)=4$.
@@
Let $F$ be differentially $4$-uniform. Then, either $\operatorname{DDT}_{F}(c,d)=2$ or $\operatorname{DDT}_{F}(c,d)=4$. One can now proof the result directly from Corollary 5.4. ∎
We will here use Corollaries 7.1– 7.2 to compute the concrete values for three infinite classes of differentially 4-uniform power permutations over $\mathbb{F}_{2^{n}}$ (Table  1) having the best known nonlinearity. For the Gold function, we will state and prove the result for general parameters, using Theorem 5.1, which implies the case for 4-differential uniform.
@@
Let $F_{1}(X)=X^{2^{s}+1}$, the general Gold function. Let $t=\gcd(s,n)$, and $m=\frac{n}{t}$. Then,
$\operatorname{EBCT}_{F_{1}}(a,b,c,d)$ =$=\begin{cases}2^{t}&\mbox{ if }{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }a=c,b=d,\mbox{ or }\\ &\quad{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }a=uc,\mbox{ and }b=(u+u^{2})c^{2^{s}+1}+ud,\\ &\qquad\qquad\mbox{ where }u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\},\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{LBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}2^{t}&\mbox{ if }{\rm Tr}_{s}^{sm}\left(\frac{c}{b^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }a\in b\mathbb{F}_{2^{t}}^{*},\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}2^{t}&\mbox{ if }{\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),c\in\{b,(u+u^{2})a^{2^{s}+1}+ub,u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\},\\ 0&\mbox{otherwise.}{}\end{cases}$
@@
(7.1) $\operatorname{EBCT}_{F_{1}}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F_{1}}(c,d)&\mbox{if}~{}\operatorname{DDT}_{F_{1}}(c,d)\neq 0,a=c\mbox{ and }b=d\\ 4\ell&\mbox{if }a\neq c,\operatorname{DDT}_{F_{1}}(c,d)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise},\end{cases}$
@@
It is known from [2, Lemma 4] and [9] that $X^{2^{s}+1}+(X+c)^{2^{s}+1}=d$ has $2^{t}$ distinct solutions if and only if $\sum_{i=0}^{m-1}D^{2^{si}}=m\equiv{\rm Tr}_{1}^{m}(1)$ (mod 2), where $D=\frac{d}{c^{2^{s}+1}}$, with $t=\gcd(s,n)$ and $m=\frac{n}{t}$, and zero solutions otherwise. Moreover, the solutions are of the form $X+cu$, where $u\in\mathbb{F}_{2^{t}}$. If $\operatorname{DDT}_{F_{1}}(c,d)=2$, by Theorem 5.1, we have $\operatorname{EBCT}_{F_{1}}(a,b,c,d)=2$ when $a=c,b=d$ and otherwise, 00.
If $t=1$, we get that $F_{1}$ is APN, and using Corollary 6.1, we are done. Next, let us suppose that $t>1$. Let $a\neq c$, then we can write $\operatorname{DDT}_{F_{1}}(c,d)=2k=2^{t}=4r$, where $r$ is a positive integer. Let $z_{i_{1}}=X+u_{i_{1}}c$ and $z_{j_{1}}=X+u_{j_{1}}c$, where $u_{i_{1}}\neq u_{j_{1}}\in\mathbb{F}_{2^{t}}$ and $u_{i_{1}}+u_{j_{1}}\neq c$. Then, for $(a,b)=(z_{i_{1}}+z_{j_{1}},F_{1}(z_{i_{1}})+F_{1}(z_{j_{1}}))=((u_{i_{1}}+u_{j_{1}})c,F_{1}(X+u_{i_{1}})+F_{1}(X+u_{j_{1}}))$, System (7.1) will have at least four solutions, namely, $\{X+u_{i_{1}}c,X+u_{j_{1}}c,X+u_{i_{1}}c+c,X+u_{j_{1}}c+c\}$, where $X$ satisfies $X^{2^{s}+1}+(X+c)^{2^{s}+1}=d$.

### v3

@@ 5. EBCT, LBCT and UBCT of a function in terms of its DDT
However, $\operatorname{DDT}_{F}(b,c)=2$ and $a\neq b$, which gives that $\operatorname{LBCT}_{F}(a,b,c)=0$. Moreover, $\operatorname{DDT}_{F}(a,b)=10$, and $S(a,b)=\{x_{1}=g,x_{2}=g^{10},x_{3}=g^{13},x_{4}=g^{28},x_{5}=g^{55},x_{1}+a=g^{19},x_{2}+a=g^{46},x_{3}+a=g^{34},x_{4}+a=g^{37},x_{5}+a=0\}$. Then, one can check that $c\neq b$ and furthermore, $c\not\in W(i,j)$ for all $i\neq j\in\{1,2,3,4\}$, implying that $\operatorname{UBCT}_{F}(a,b,c)=0$.
@@
Let $F$ be a function on $\mathbb{F}_{2^{n}}$. Given any $a,b,c,d\in\mathbb{F}_{2^{n}}$, the following statements are equivalent:
@@
On the other hand (replacing the input $(a,b,c)$ by $(a,c,d)$),
$\operatorname{LBCT}_{F}(a,c,d)=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists~{}Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}(4)\ X+Y=c\\ (5)\ F(X+a)+F(Y+a)=d\\ (6)\ F(X)+F(Y)=d\end{cases}\right\}\right\rvert.$
@@
Furthermore, (replacing the input $(a,b,c)$ by $(c,d,b)$),
$\operatorname{UBCT}_{F}(c,d,b)=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists~{}Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}(7)\ F(X+c)+F(Y+c)=b\\ (8)\ F(X)+F(Y)=b\\ (9)\ F(X)+F(X+c)=d\end{cases}\right\}\right\rvert.$
@@ 6. Consequences of our results for APN functions
If $F$ is an APN function over $\mathbb{F}_{2^{n}}$, then for $a,b,c,d\in\mathbb{F}_{2^{n}}$, we have
$\operatorname{EBCT}_{F}(a,b,c,d)$ =$=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if}~{}c=d=0,\\ \operatorname{DDT}_{F}(c,d)&\mbox{if}~{}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{ or }a=0,b=0\mbox{ and }c\neq 0,\\ 0&\mbox{otherwise},\end{cases}$
$\operatorname{LBCT}_{F}(a,b,c)$ =$=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}b=c=0;\mbox{ or }a=b\mbox{ and }ab\neq 0;\mbox{ or }a=0\mbox{ and }b\neq 0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F}(a,b,c)$ =$=\begin{cases}|F^{-1}(c+{\rm Im}(F))|&\mbox{if}~{}a=0\mbox{ and }b=0,\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}c=b\mbox{ and }a\neq 0,\\ 0&\mbox{otherwise,}\end{cases}$
where ${\rm Im}(F)$ is the image of $F$ and $F^{-1}(\cdot)$ denotes the preimage of the argument.
@@
For any $a,b,c,d\in\mathbb{F}_{2^{n}}$, $(\operatorname{EBCT}_{F}(a,b,c,d))^{2}\leq\operatorname{LBCT}_{F}(a,c,d)\cdot\operatorname{UBCT}_{F}(c,d,b)$. Furthermore, the equality is met for every $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$ if and only if $F$ is APN.
@@
The inequality follows directly from Theorem 5.3. Suppose that $F$ is APN. Then, by Corollary 6.1, $(\operatorname{EBCT}_{F}(a,b,c,d))^{2}=\operatorname{LBCT}_{F}(a,c,d)\cdot\operatorname{UBCT}_{F}(c,d,b)$. Suppose now that $F$ is not APN. Then, for some $c,d\in\mathbb{F}_{2^{n}}^{*}$, there exist at least four distinct solutions of the equation $F(X+c)+F(X)=d$, which we denote by $z_{1},z_{2},z_{1}+c,z_{2}+c$. Let now $a=z_{1}+z_{2},b=F(z_{1})+F(z_{2})+d$. Then, since $a\neq c$ (otherwise $z_{2}=z_{1}+c$), $(\operatorname{EBCT}_{F}(a,b,c,d))^{2}<\operatorname{LBCT}_{F}(a,c,d)\cdot\operatorname{UBCT}_{F}(c,d,b)$. ∎
@@ 7. Consequences of our results for some well-known functions
In this section, we will give the EBCT, LBCT and UBCT entries of the Gold, Kasami and Bracken-Leander functions for those parameters such that the functions are differentially 4-uniform permutations; in the case of the Gold function, we also state and prove the EBCT, LBCT and UBCT entries for any parameters. We can also give a very short proof for the EBCT, LBCT and UBCT entries of the inverse function for $n$ even, which is the main result in [13, 17] (for $n$ odd, see the previous section).
For clarity, we first state the results we will use to compute the EBCT, LBCT and UBCT entries of differentially $4$-uniform functions using Theorem 5.1 and Theorem 5.3, which we will need for the functions in this section.
@@
Let $F$ be a differentially $4$-uniform function over $\mathbb{F}_{2^{n}}$. Then, for $a,b,c,d\in\mathbb{F}_{2^{n}}$, we have
$\operatorname{EBCT}_{F}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if }c=d=0,\\ \operatorname{DDT}_{F}(c,d)&\mbox{if}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{ or }a=0,b=0\mbox{ and }c\neq 0;\mbox{ or }\\ &~{}\operatorname{DDT}_{F}(c,d)=4,a=z_{1}+z_{2}\mbox{ and }b=F(z_{1})+F(z_{2});\mbox{ or }\\ &~{}\operatorname{DDT}_{F}(c,d)=4,a=z_{1}+z_{2}+c\mbox{ and }b=F(z_{2})+F(z_{2}+c),\\ 0&\mbox{otherwise},\end{cases}$
where $z_{1},z_{1}+c,z_{2},z_{2}+c$ are the four solutions of the equation $F(X+c)+F(X)=d$, if $\operatorname{DDT}_{F}(c,d)=4$. Next,
$\operatorname{LBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}b=c=0;\mbox{ or }a=0\mbox{ and }b\neq 0;\mbox{ or }\\ &~{}\operatorname{DDT}_{F}(b,c)=2,a=b\mbox{ and }ab\neq 0;\mbox{ or }\\ &~{}\operatorname{DDT}_{F}(b,c)=4,ab\neq 0\mbox{ and }a\in\{b,y_{1}+y_{2},y_{1}+y_{2}+b\},\\ 0&\mbox{otherwise},\end{cases}$
where $y_{1},y_{1}+b,y_{2},y_{2}+b$ are the four solutions of the equation $F(X+b)+F(X)=c$, if $\operatorname{DDT}_{F}(b,c)=4$. Further,
$\operatorname{UBCT}_{F}(a,b,c)=\begin{cases}|F^{-1}(c+{\rm Im}(F))|&\mbox{if}~{}a=b=0,\\ \operatorname{DDT}_{F}(a,b)&~{}\operatorname{DDT}_{F}(a,b)=2,a\neq 0\mbox{ and}c=b;\mbox{ or }\\ &~{}\operatorname{DDT}_{F}(a,b)=4,a\neq 0\mbox{ and }\\ &~{}c\in\{b,F(x_{1})+F(x_{2}),F(x_{1})+F(x_{2})+b\},\\ 0&\mbox{otherwise},\end{cases}$
where $x_{1},x_{1}+a,x_{2},x_{2}+a$ are the four solutions of the equation $F(X+a)+F(X)=b$, if $\operatorname{DDT}_{F}(a,b)=4$, ${\rm Im}(F)$ is the image of $F$ and $F^{-1}(\cdot)$ denotes the preimage of the argument.
@@
Here, we give an alternative characterization of EBCT of a differentially $4$-uniform function $F$ using LBCT and UBCT of $F$.
@@
Let $F$ be differentially $4$-uniform. Then, since $cd\neq 0$, either $\operatorname{DDT}_{F}(c,d)=0$, $\operatorname{DDT}_{F}(c,d)=2$ or $\operatorname{DDT}_{F}(c,d)=4$. One can now proof the result directly from Theorem 5.3. ∎
We will here use Corollaries 7.1– 7.2 to compute the concrete values for three infinite classes of differentially 4-uniform power permutations over $\mathbb{F}_{2^{n}}$ (Table  1) having the best known nonlinearity. For the Gold function, using Theorem 5.1, we will state and prove the result for general parameters, including those for which the Gold function is 4-differential uniform.
@@
Let $F_{1}(X)=X^{2^{s}+1}$ be the Gold function on $\mathbb{F}_{2^{n}},1\leq s<n$. Let $t=\gcd(s,n)$, and $m=\frac{n}{t}$. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}$, we have
$\operatorname{EBCT}_{F_{1}}(a,b,c,d)$ =$=\begin{cases}2^{n}&\mbox{if }a=b=c=d=0,\\ 2^{t}&\mbox{if }{\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),c=d=0\mbox{ and }a\neq 0;\mbox{ or }\\ &~{}{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),a=b=0\mbox{ and }c\neq 0;\mbox{ or }\\ &~{}{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),a=c,b=d\mbox{ and }ac\neq 0;\mbox{ or }\\ &~{}{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),a=uc,ac\neq 0\mbox{ and }\\ &~{}b=(u+u^{2})c^{2^{s}+1}+ud,\mbox{ where }u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\},\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{LBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}2^{n}&\mbox{if }b=c=0,\\ 2^{t}&\mbox{if }{\rm Tr}_{s}^{sm}\left(\frac{c}{b^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),a=0\mbox{ and }b\neq 0;\mbox{ or }\\ &~{}{\rm Tr}_{s}^{sm}\left(\frac{c}{b^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1),a\in b\mathbb{F}_{2^{t}}^{*}\mbox{ and }ab\neq 0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}|F_{1}^{-1}(c+{\rm Im}(F_{1}))|&\mbox{if }~{}a=b=0,\\ 2^{t}&\mbox{if }a\neq 0,{\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }\\ &~{}c\in\{b,(u+u^{2})a^{2^{s}+1}+ub,u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\},\\ 0&\mbox{otherwise,}{}\end{cases}$
where ${\rm Im}(F_{1})$ is the image of $F_{1}$ and $F_{1}^{-1}(\cdot)$ denotes the preimage of the argument. In particular when $F_{1}$ is a permutation or equivalently when $m$ is odd, then $|F_{1}^{-1}(c+{\rm Im}(F_{1}))|=2^{n}$.
@@
$\operatorname{EBCT}_{F_{1}}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F_{1}}(a,b)&\mbox{if }c=d=0,\\ \operatorname{DDT}_{F_{1}}(c,d)&\mbox{if}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{or }a=b=0\mbox{ and }c\neq 0,\\ 4\ell&\mbox{if }a\neq c,ac\neq 0\mbox{ and }\operatorname{DDT}_{F_{1}}(c,d)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise},\end{cases}$
@@
It is known from [2, Lemma 4] and [10] that $X^{2^{s}+1}+(X+c)^{2^{s}+1}=d$ has $2^{t}$ distinct solutions if and only if $\sum_{i=0}^{m-1}D^{2^{si}}=m\equiv{\rm Tr}_{1}^{m}(1)$ (mod 2), where $D=\frac{d}{c^{2^{s}+1}}$, with $t=\gcd(s,n)$ and $m=\frac{n}{t}$, and zero solutions otherwise. Moreover, the solutions are of the form $X+cu$, where $u\in\mathbb{F}_{2^{t}}$.
If $t=1$, we get that $F_{1}$ is APN, and using Corollary 6.1, we are done. Next, let us suppose that $t>1$. Here, we focus exclusively on the case where $a\neq c$ and $ac\neq 0$, as all other cases are straightforward. Let $\operatorname{DDT}_{F_{1}}(c,d)=2k=2^{t}=4r$, where $r$ is a positive integer. Let $z_{i_{1}}=X+u_{i_{1}}c$ and $z_{j_{1}}=X+u_{j_{1}}c$, where $u_{i_{1}}\neq u_{j_{1}}\in\mathbb{F}_{2^{t}}$ and $u_{i_{1}}+u_{j_{1}}\neq c$. Then, for $(a,b)=(z_{i_{1}}+z_{j_{1}},F_{1}(z_{i_{1}})+F_{1}(z_{j_{1}}))=((u_{i_{1}}+u_{j_{1}})c,F_{1}(X+u_{i_{1}})+F_{1}(X+u_{j_{1}}))$, System (3.1) will have at least four solutions, namely, $\{X+u_{i_{1}}c,X+u_{j_{1}}c,X+u_{i_{1}}c+c,X+u_{j_{1}}c+c\}$, where $X$ satisfies $X^{2^{s}+1}+(X+c)^{2^{s}+1}=d$.